OpenStudy (anonymous):

Simplify 18y^3/36xy^2

OpenStudy (anonymous):

1/2x y/2x y^2/2x 1/2xy

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

\[\frac{ 18y^3 }{ 36xy^2 } \\= \frac{ 18*1*y*y*y }{ 18*2*x*y*y} \\= \\\\ \frac{ y }{2x }\]